OpenStudy (alprincenofl):

The recursive definition of the set S of positive integers that are multiples of 5 is

OpenStudy (alprincenofl):

http://prntscr.com/57sph9

OpenStudy (anonymous):

Consider letting \(x = 5n\) and \(y=5m\), then ensuring all of them would still be multiples of \(5\).

OpenStudy (anonymous):

For example \[ x+y = 5n+5m = 5(m+n) \]So it's still multiple of \(5\). The next step is to figure out whether we cover all multiples of \(5\).

OpenStudy (anonymous):

For example, consider the second option: \[ 5y = 5(5n) = 5^2n \]We will really only get numbers of the form \(5^kn\). When we say \(5\in S\) it let's us know that \(n=1\) to start with. We can't get \(10 = 2\times 5\) because there is no \(k\) where \(5^kn = 10\). So the second option won't work because it doesn't include all multiples of 5.

OpenStudy (alprincenofl):

Thank you, Much appreciated!